# Evaluating World Models with LLM for Decision Making

### v1

@@ Abstract
World model emerges as a key module in decision making, where MuZero and Dreamer achieve remarkable successes in complex tasks. Recent work leverages Large Language Models (LLMs) as general world simulators to simulate the dynamics of the world due to their generalizability. LLMs also serve as the world model for deliberative reasoning in Reasoning via Planning (RAP) and Tree of Thought (ToT). However, the world models are either evaluated as a general world simulator, or as a functional module of the agent, i.e., predicting the transitions to assist the planning. In this work, we propose a comprehensive evaluation of the world models with LLMs from the decision making perspective. Specifically, we leverage the 31 diverse environments from (Wang et al., 2023; 2024) and curate the rule-based policy of each environment for the diverse evaluation. Then, we design three main tasks, i.e., policy verification, action proposal, and policy planning, where the world models can be used for decision making solely. Finally, we conduct the comprehensive evaluation of the advanced LLMs, i.e., GPT-4o and GPT-4o-mini, on the environments for the three main tasks under various settings. The key observations include: i) GPT-4o significantly outperforms GPT-4o-mini on the three main tasks, especially for the tasks which require the domain knowledge, ii) the performance of the world model with LLM will be decreased for long-term decision-making tasks, and iii) the combination of different functionalities of the world model will brings additional unstabilities of the performance.
@@ 1 Introduction
Due to the celebrating success of MuZero (Schrittwieser et al., 2020) and Dreamer (Hafner et al., 2019; 2021; 2023), world model (Ha & Schmidhuber, 2018) becomes a key concept in decision making. With the encoding vast knowledge of the world for the prediction of the new states after taking the actions, world models have demonstrate their effectiveness for generalizing to novel tasks (Byravan et al., 2020), efficient planning with world model (Sekar et al., 2020), and working on offline datasets (Schrittwieser et al., 2021; Yu et al., 2020; 2021). World models can also be viewed as general world simulators where users can interact with, e.g., Genie (Bruce et al., 2024) and Vista (Gao et al., 2024).
Large Language Models (LLMs) achieve remarkable success in enormous natural language tasks in the past five years (Brown et al., 2020; OpenAI, 2023). Several recent works leverage LLMs as the generalist world models to provide the environment knowledge for various complex tasks, e.g., math and reasoning. With the fine-tuning over pre-collected data from the environments, the LLMs can predict the action sequences across different tasks over environments while maintaining the capabilities on other domains (Xiang et al., 2023). LLMs also serve as the world model explicitly in Reasoning via Planning (RAP) (Hao et al., 2023) and Reason for Future, Act for Now (RAFA) (Liu et al., 2023), where the LLMs predict the next states based on the actions executed at current states, e.g., the states of blocks in the BlocksWorld (Valmeekam et al., 2023), which is used to assist the planning methods. On the other hand, LLMs serve as the world model implicitly in the widely-used Tree of Thoughts (ToT) (Yao et al., 2023), as well as Graph of Thoughts (GoT) (Besta et al., 2024), where the LLMs need to predict the states and evaluate the thoughts. Recent work also consider LLMs as world simulators (Wang et al., 2024; Xie et al., 2024), where they evaluate the performance of LLMs on the prediction of next states and the game progress, demonstrating the potentials of LLMs as general world models.
Figure 1: Picking the action with higher value. The prediction of B in Case 2 is more accurate than Case 1 in term of the $L_{2}$ loss, but leads to the wrong action.
However, most of the previous works evaluate the world models with LLMs either as general world simulators (Wang et al., 2024), or as additional modules of the agents to make decisions (Liu et al., 2023). We argue that a comprehensive evaluation of the world models themselves from the decision-making perspective is needed due to the two facts. First, the objective of decision making is finding the policy to complete the task, where only a small portion of the world will be visited during finding and executing the policy, given the fact that AlphaZero can find the super-human policy (Silver et al., 2018) by only exploring a small proportion (less than 1%) of the state space, therefore, the evaluation of the world models on the predictions of the transitions that relevant to the desired policy is more important than the transitions far from the policy. Second, the influence of the world models is usually coupled with the actors who choose the actions, i.e., if the actor cannot pick the correct actions, the task cannot be completed even when the world model is accurate, which brings additional difficulties to understand the performances of the world models.
Therefore, we address these issues with three key observations in this work:
  • Observation 1: Prediction is important, but not that important. An illustrative example is displayed in Figure 1, which indicates that more accurate predictions do not lead to right decisions. [^1] The issue in Figure 1 can be elicited by various methods, e.g., rank prediction. This example is just to illustrate the discrepancy between prediction and decision, motivating us to reconsider the evaluation of the world model for decision making. This phenomenon is also observed in other decision making scenarios, e.g., financial trading (Sun et al., 2023). The success of MuZero Unplugged (Schrittwieser et al., 2021) also demonstrate that we can learn good policies from inaccurate world models which are trained only with limited data. This motivates us that the evaluation of the world models for decision making should focus on the predictions which relevant to the desired policy, rather than as general world simulators.
  • Observation 2: Selecting potential actions should be an important feature for world models. Most of the previous works in world model focus on next state and reward prediction, and the action selection is usually completed by separate model. Several works (Xie et al., 2024; Janner et al., 2021) incorporate the action selection into consideration. We argue that as the world model for decision making has more knowledge about the world, which can potentially make a better selection of the potential actions. World models can also be viewed as game engines (Valevski et al., 2024), which have to provide potential actions to guide fresh players to complete tasks, e.g., Red Dead Redemption 2 (Tan et al., 2024).
  • Observation 3: Planning with world models can find the policies solely. With the prediction of the next states and the action proposal, we can leverage planning methods or search methods to find the policies. Most works introduce the critic (i.e., the value function) to evaluate the actions immediately for efficient planning (Schrittwieser et al., 2020; Hao et al., 2023), however, the critic is not necessary for finding policies and may also influence the performance. Therefore, we should evaluate the world model solely for finding the policies to avoid the influences of other modules.
Figure 2: Evaluation of World Model with LLM for Decision Making.

### v2

# But Rigorous Evaluations are Needed

@@ Abstract
World model emerges as a key module in decision making, where MuZero and Dreamer achieve remarkable successes in complex tasks. Recent work leverages Large Language Models (LLMs) as general world simulators to simulate the dynamics of the world due to their generalizability. LLMs also serve as the world model for deliberative reasoning in Reasoning via Planning (RAP) and Tree of Thought (ToT). However, the world model is either evaluated as a general world simulator, or as a functional module of the agent, i.e., predicting the transitions to assist the planning. This paper argues that LLM-based world models can make decisions solely, but rigorous evaluations are needed. We first present the two key observations to showcase how LLM-based world models can make decisions solely, and then present the three key observations to demonstrate why current evaluation framework of LLM-based world models is not sufficient. Then, we present our suggested evaluation framework: policy verification, action proposal, and policy planning, where the world model is used for decision making solely, and finally we leverage the 31 diverse environments from (Wang et al., 2023; 2024) and curate the rule-based policy of each environment for diverse evaluations. The key findings include: i) GPT-4o significantly outperforms GPT-4o-mini on the three main tasks, especially for the tasks which require the domain knowledge, e.g., scientific tasks, ii) the performance of the LLM-based world models depends predominantly on their performance in key steps, while the total number of steps required for task completion is a weak indicator of task difficulty with critical bottleneck steps playing a more decisive role and iii) the combination of world models’ functionalities for decision making tends to increase performance instability in our experiments, which can partially obscure the performance gap between stronger and weaker model.
@@ 1 Introduction
The remarkable achievements of MuZero (Schrittwieser et al., 2020) and Dreamer (Hafner et al., 2019; 2021; 2025a; 2025b) have established world models (Ha and Schmidhuber, 2018) as fundamental components in decision-making systems. World models serve as learned simulators that encode rich representations of environment dynamics, enabling agents to predict future states conditioned on their actions. Rather than learning direct policy mappings from observations to actions, world models learn the underlying transition dynamics $T(s,a)\rightarrow s^{\prime}$ and reward functions $R(s,a)$, providing a compressed yet informative representation of how the environment evolves. These models have demonstrated effectiveness across several key areas with concrete achievements: i) Generalization to novel tasks. World models enable transfer learning by capturing reusable environment dynamics (Byravan et al., 2020). For instance, agents trained with world models can adapt to new task objectives without relearning dynamics (Robey et al., 2021), and demonstrate robust performance across distribution shifts in robotics applications (Young et al., 2023). This capability is particularly valuable when the cost of re-training from scratch is prohibitive. ii) Efficient planning. By simulating possible futures in the learned model, agents can evaluate action sequences through "mental simulation" rather than costly real-world interaction (Sekar et al., 2020; Hamrick et al., 2021). MuZero leverages this capability to achieve superhuman performance in Go, chess, and Atari games by planning over imagined trajectories (Schrittwieser et al., 2020). The ability to explore hypothetical scenarios without environmental interaction dramatically improves sample efficiency and enables safer exploration in high-risk domains. iii) Offline learning from fixed datasets. World models have proven especially valuable in offline reinforcement learning settings (Schrittwieser et al., 2021; Yu et al., 2020; 2021), where agents must learn purely from pre-collected datasets without additional environment interaction. By learning accurate dynamics models from historical data, agents can perform policy optimization in the model space, generating synthetic rollouts to augment limited real data. This capability opens possibilities for training agents in domains where online interaction is impractical or expensive, such as healthcare and autonomous driving.
Recent developments have extended world models beyond traditional reinforcement learning to interactive simulation systems, demonstrating their potential as general-purpose world understanding engines. Notable examples include the Genie series (Bruce et al., 2024; Parker-Holder et al., 2024; Ball et al., 2025), which learns controllable world models from unlabeled video data and enables users to interact with generated game-like environments through action inputs. Vista (Gao et al., 2024) extends this paradigm to autonomous driving, learning world models that can simulate diverse driving scenarios and enable policy training in purely synthetic environments. Additional systems such as MatrixGame (Zhang et al., 2025; He et al., 2025), which generates playable game environments from natural language descriptions, and Hunyuan-GameCraft (Li et al., 2025), which creates interactive 3D game worlds, further illustrate the expanding capabilities of world models in interactive domains. These systems share a common theme: they learn predictive models of complex, high-dimensional dynamics (visual observations, physics, agent interactions) and expose these models through interactive interfaces. These advances point toward world models serving as foundational components for artificial general intelligence systems with grounded understanding of world dynamics, moving beyond narrow task-specific applications to general-purpose simulation and interaction.
Large Language Models (LLMs) have emerged as a promising approach to world modeling, leveraging their broad knowledge acquired from large-scale text corpora. Building on their success in natural language tasks (Brown et al., 2020; OpenAI, 2023), recent works explore LLMs as general world models that provide environment knowledge for complex reasoning and decision-making tasks. The key hypothesis is that LLMs, having been exposed to vast amounts of text describing how the world works, have internalized implicit models of common-sense physics, social dynamics, and domain-specific knowledge (e.g., mathematics, science, games). Through fine-tuning on environment interaction data, i.e., sequences of states, actions, and outcomes, LLMs can predict action sequences across diverse tasks while maintaining their broad capabilities (Xiang et al., 2023; Wang et al., 2024; Xie et al., 2024). For example, an LLM fine-tuned on household task demonstrations can predict valid action sequences for cooking or cleaning tasks while retaining its language understanding and reasoning abilities. LLMs serve as world models in both explicit and implicit ways, depending on how the state prediction capability is utilized: i) Explicit world modeling. In approaches like Reasoning via Planning (RAP) (Hao et al., 2023) and Reason for Future, Act for Now (RAFA) (Liu et al., 2023), LLMs directly predict next states given current states and actions, functioning as explicit transition models $T_{\text{LLM}}(s_{t},a_{t})\rightarrow s_{t+1}$. For instance, in BlocksWorld (Valmeekam et al., 2023), given a state description like “Block A is on Block B, Block C is on the table” and an action “move Block A to Block C,” the LLM predicts the resulting state: “Block A is on Block C, Block B is on the table, Block C is on the table.” These predicted states are then used by planning algorithms (e.g., Monte Carlo Tree Search, breadth-first search) to evaluate action sequences and select optimal plans. The explicit separation allows for modular analysis of world model accuracy independent of planning algorithm performance. ii) Implicit world modeling. In widely-used approaches like Tree of Thoughts (ToT) (Yao et al., 2023) and Graph of Thoughts (GoT) (Besta et al., 2024), LLMs perform world modeling implicitly as part of their reasoning process. Rather than explicitly outputting state descriptions, these methods have LLMs generate and evaluate intermediate reasoning steps (“thoughts”) that implicitly represent states in a problem-solving trajectory. For example, when solving a math problem, each thought represents a partial solution state, and the LLM must implicitly model how different reasoning steps transition between these states. The LLM evaluates the promise of each thought (analogous to value estimation in RL) to guide search through the solution space. We provide a comprehensive review of related work in Section 6.